OpenStudy (anonymous):

x^-1+y^-1 over x+y eliminate the complex fractions

OpenStudy (anonymous):

\[\frac{x^{-1}+y^{-1}}{x+y}=\frac{\frac{1}{x}+\frac{1}{y}}{x+y}\] clear the compound fraction by multiplying top and bottom by \(xy\) as a first step